# Online Energy Storage Arbitrage under Imperfect Predictions: A Conformal Risk-Aware Approach

### v1

@@ Appendix B Analytical Solution to Energy Storage Arbitrage
The following results from (Zheng et al., 2022) are reproduced here for reader completeness and to facilitate interpretation of our risk-aware dispatch policy (13).
@@
Appendix C Proof of Theorem 2
@@
Appendix D Proof of Proposition 2
@@
According to Remark 2, a dispatch policy that can maximize the total profit maintains an optimal total value constantly at any time step $t$, i.e., $V_{t}:=\Pi_{t}(p^{\prime}_{t},b^{\prime}_{t}|\lambda_{t})+Q_{t}(e^{\prime}_{t+1})$, where $\Pi_{t}$ and $e^{\prime}_{t+1}$ are updated with:
@@
Appendix E Proof of Proposition 5
@@
Following Remark 4, we can state that a conservative policy is less sensitive to changes in forecast uncertainty, and vice versa. That is, for any two value functions $Q_{1},Q_{2}$ and $0<\gamma_{1}<\gamma_{2}$, the following holds:
@@
Appendix F Baseline Algorithms
@@
    Chance Constrained: chance-constrained optimization ensuring constraint violations occur with probability below specified thresholds $\Gamma$, but requiring accurate probability distributions and lacks adaptability to distributional shifts (Wu et al., 2025). Let $\xi$ be an epigraph auxiliary variable:
@@
    Robust Optimization: feasible solutions guarantee under worst-case realizations within predefined uncertainty sets $\mathcal{U}_{t}$, but tends conservative and cannot adapt uncertainty set specifications online (Wu et al., 2025):
@@
    Switching Cost: an online prediction-based algorithm conversion with switching costs to handle noisy predictions (Chen et al., 2015), though its theoretical risk management guarantees assume observable prediction errors, which are not available in our setting. Let $\zeta$ be the switching cost penalty factor:

### v2

@@ Appendix B Analytical Solution to Energy Storage Arbitrage
The following results from (zheng2022) are reproduced here for reader completeness and to facilitate interpretation of our risk-aware dispatch policy (14).
@@
Appendix C Proof of Proposition 3
Proof.
$\Delta(e_{t})$ $=\hat{Q}_{t}(e_{t}|\lambda_{t})-Q_{t}(e_{t})$
$=\hat{Q}_{t}(e_{t}|\lambda_{t})-(\Pi_{t}+Q_{t}(e_{t+1}))$
$(\Rightarrow)$ First we show that if $\delta(e_{t})=0$, then $\Delta(e_{t})=0$, for all $e_{t}\in[0,E],\ t\in\mathcal{T}$. Given the recursive relationship established above, for a given time step $t$, $\Delta(e_{t})=\delta(e_{t})+\Delta(e_{t+1})=0$ if $\Delta(e_{t+1})=0$ for all $e_{t+1}\in[0,E]$, since $\delta(e_{t})=0$. With the boundary condition $\Delta(e_{T})=0$ for all $e_{T}\in[0,E]$ at $t=T$, it follows by induction that $\Delta(e_{t})=0$ if $\delta(e_{t})=0$ for all $e_{t}\in[0,E],\ t\in\mathcal{T}$.
Appendix D Proof of Theorem 2
@@
Appendix E Proof of Proposition 4
@@
According to Section 3.4, a dispatch policy that can maximize the total profit maintains an optimal total value constantly at any time step $t$, i.e., $V_{t}:=\Pi_{t}(p^{\prime}_{t},b^{\prime}_{t}|\lambda_{t})+Q_{t}(e^{\prime}_{t+1})$, where $\Pi_{t}$ and $e^{\prime}_{t+1}$ are updated with:
@@
Appendix F Proof of Proposition 4
@@
Following Proposition 4, we can state that a conservative policy is less sensitive to changes in forecast uncertainty, and vice versa. That is, for any two value functions $Q_{1},Q_{2}$ and $0<\gamma_{1}<\gamma_{2}$, the following holds:
@@
Appendix G Baseline Algorithms
@@
    Chance Constrained: chance-constrained optimization ensuring constraint violations occur with probability below specified thresholds $\Gamma$, but requiring accurate probability distributions and lacks adaptability to distributional shifts (wu2025). Let $\xi$ be an epigraph auxiliary variable:
@@
    Robust Optimization: feasible solutions guarantee under worst-case realizations within predefined uncertainty sets $\mathcal{U}_{t}$, but tends conservative and cannot adapt uncertainty set specifications online (wu2025):
@@
    Switching Cost: an online prediction-based algorithm conversion with switching costs to handle noisy predictions (chen2015), though its theoretical risk management guarantees assume observable prediction errors, which are not available in our setting. Let $\zeta$ be the switching cost penalty factor:
@@
Appendix H Conformal Controller Sensitivity Analysis
We include here an extended sensitivity analysis of key algorithm parameters, including the controller learning rate $\rho$ in Eq. (11) (Figure 7) and the sensitivity parameter $\sigma$ in Eq. (13) (Figure 8). The layout of each figure is similar to that of Figure 6.
H.1. Impact of Controller Learning Rate
We analyze the sensitivity of the proposed algorithm to the conformal controller calibration learning-rate parameter $\rho$ in Eq. (11), which governs how quickly the control variable adjusts in response to observed losses.
Under good forecast quality (Figure 7(7(a))), both CC-prediction err and CC-value err exhibit relatively mild variation across the tested range of $\rho$. This robustness arises because accurate forecasts stabilize the temporal-difference signal that drives calibration updates. Note that CC-value err shows slightly higher sensitivity to $\rho$ than CC-prediction err, as the the prediction-error-based updates leverage the heuristic conservativeness calibration function $f$, which helps guide the controller toward more stable convergence.
Under poor forecast quality (Figure 7(7(b))), larger learning rates cause the controller to update too aggressively. As $\rho$ increases, both the cumulative risk $\hat{R}_{T}$ and realized profit loss increase significantly for both the prediction-error-based and value-error-based calibration methods. This indicates that when forecasts are noisy, overly large $\rho$ magnifies volatility in decision conservativeness and exposes the system to greater downside risk.
H.2. Impact of Prediction Set Scaling Parameter
We conduct a sensitivity analysis to examine how the conformal controller responds to variations in the prediction set scaling parameter $\sigma$ in Eq. (13) under different levels of forecast accuracy. A larger $\sigma$ produces wider prediction sets and therefore more conservative dispatch decisions, whereas a smaller $\sigma$ yields narrower prediction sets and more aggressive actions driven by higher confidence in the forecasted opportunity value.
Under good forecast quality (Figure 8(8(a))), the results are similar to that in Figure 7(7(a)).
Under poor forecast quality (Figure 8(8(b))), an overly narrow prediction set (small $\sigma$) leads to significant increases in both cumulative risk $\hat{R}_{T}$ and realized profit loss, as the controller becomes overconfident and vulnerable to prediction errors, violating the conditions of Proposition 4. Conversely, given overly wide prediction sets (large $\sigma$), the cumulative risk $\hat{R}_{T}$ remains low, but the realized profit loss for CC-prediction-err increases. This effect is likely due to inappropriate estimate of heuristic conservativeness calibration function $f$.
(a) $R^{2}=0.4$.
(b) $R^{2}=-0.4$.
Figure 7. Impact of learning rate $\rho$ considering good-accuracy ($R^{2}=0.4$) and poor-accuracy ($R^{2}=-0.4$) forecasters.
(a) $R^{2}=0.4$.
(b) $R^{2}=-0.4$.
Figure 8. Impact of prediction set sensitivity parameter $\sigma$ considering good-accuracy ($R^{2}=0.4$) and poor-accuracy ($R^{2}=-0.4$) forecasters.